OpenStudy (anonymous):

11s-5=13s+13 Answer please.

OpenStudy (anonymous):

s=-9

OpenStudy (anonymous):

Re-arranging the equation by moving s on left and rest values on right, we get: 11s-13s=13+5 -2s=18 s=-9

OpenStudy (anonymous):

@Shyguy77 this one too 10y-12=11y-7

OpenStudy (anonymous):

y=-5

OpenStudy (anonymous):

@Shyguy77 -7=1/2x+5

OpenStudy (anonymous):

-7*(2x+5)=1 -14x-35=1 -14x=36 x=-36/14 x=-18/7

OpenStudy (anonymous):

the 1/2 is a fraction

OpenStudy (anonymous):